Solve the system of equations by graphing:
2x-6y=24
3x-9y=-54

Answer by mananth(16946)

equation1 #DIV/0! = 24
when x= 0 y= -4 ( 0 , -4 )
when y = 0 x= 12 (12 , 0 )
Equation 2
when x= 0 y= 6 ( 0 , 6 ) 6
when y = 0 x= -18 ( -18 , 0 )